Your firm needs a computerized machine tool lathe which costs $85,452 and requires $10,263 in maintenance expense for each year of its 5-year life. After five years, this machine will be replaced. The machine falls into the MACRS (Links to an external site.) 5-year class life category. Assume a tax rate of 25 percent and a discount rate of 12 percent. If the lathe can be sold for $7,281 at the end of year 5, what is the after-tax salvage value? Below are the MACRS rates for the 5-year recovery period. DO NOT USE DOLLAR SIGNS OR COMMAS IN YOUR ANSWER. ENTER YOUR ANSWER TO THE NEAREST DOLLAR (e.g. 1250).

Year MACRS Rate
1 20.00%
2 32.00%
3 19.20%
4 11.52%
5 11.52%
6 5.76%
Answer #1

2 А в с Year Cost Depreciation Accumulated Ending book Depreciation rate expenses depreciation value 95715 20% 19143 19143 76

Excel formula:

ДА В Ε F Year 2 0 au wn Cost =85452+10263 |=85452+10263 =85452+10263 =85452+10263 =85452+10263 Depreciation rate 0.2 0.32 0.1